The measure of the supplement of an angle is 3 times thre measure of the complement of the angle. Find the measuer of the angle
Answer by lwsshak3(11628)

The measure of the supplement of an angle is 3 times thre measure of the complement of the angle. Find the measuer of the angle
**
let x=angle
supplement of angle=180-x
complement of angle=90-x
180-x=3(90-x)
180-x=270=3x
2x=90
x=45º